If the price of a $15 T-shirt is reduced by 1/3, how much sales tax must be paid on the T-shirt in a state having a 4% sales tax?
Answer by Paul(988)

1/3 = 0.33
15/3 = 5
5 x 0.33
4.95
15-4.95 = 10.05
10.05 x 0.04 = 0.40
Therefore, the sales tax must be $0.40